# Do graph neural network states contain graph properties?

### v1

@@ Abstract
Graph learning models achieve state-of-the-art performance on many tasks, but this often requires increasingly large model sizes. Accordingly, the complexity of their representations increase. Explainability techniques (XAI) have made remarkable progress in the interpretability of ML models. However, the non-relational nature of Graph Neural Networks (GNNs) make it difficult to reuse already existing XAI methods. While other works have focused on instance-based explanation methods for GNNs, very few have investigated model-based methods and, to our knowledge, none have tried to probe the embedding of the GNNs for well-known structural graph properties. In this paper we present a model agnostic explainability pipeline for Graph Neural Networks (GNNs) employing diagnostic classifiers. This pipeline aims to probe and interpret the learned representations in GNNs across various architectures and datasets, refining our understanding and trust in these models.
@@ 1 Introduction
In the last decade, significant progress has been made towards modelling non-Euclidean, graph-structured data (Kipf & Welling, 2017) on the one hand, and on interpreting the predictions of deep neural networks (DNN) on the other hand. We often qualify DNNs as black box as their predictions are not inherently interpretable. Occlusion, gradient, perturbation, layer-wise relevance propagation, and attention mechanisms have been proposed to solve this problem (Zeiler & Fergus, 2013; Denil et al., 2015; Li et al., 2016; Sundararajan et al., 2017). These methods focus on highlighting the importance of different input features. They can, however, not be directly applied on GNNs due to the lack of a regular structure (e.g. vertices can have different degrees). In this case, explaining a prediction means identifying important parts of the relational structure, or input features of nodes. An issue is that finding the explanation is itself a combinatorial problem, making XAI (explainable AI) methods for GNN intractable (Longa et al., 2023a; Ying et al., 2019; Lucic et al., 2022).
@@
Previous surveys (Agarwal et al., 2023; Dai et al., 2022) highlighted the lack of comprehensive, robust and model-agnostic explainability methods. We also identified that there are very few model-level explainability methods. As an alternative to these more traditional XAI methods, we propose to apply probing techniques for graph properties (as developed for Natural Language Processing Giulianelli et al. (2018), Belinkov (2021)) to GNN embeddings. In our pipeline (see fig. 1), we investigate both local properties like betweennes centrality, as well as global properties like average path length. To our knowledge, this is the first work to explore this direction.
@@
Figure 1: An example of the probing pipeline. First, a GNN is trained on a specific task, for example detecting whether a graph contains a grid or house shaped pattern. Then, we extract embeddings from the internal layers of the network. We use these embeddings to train the probing model; in this example a binary classifier which can detect whether the embedding contains predictive information for the diameter of the graph. If a linear probe has good performance ($R^{2}$ score) then there exists a hyperplane in the representation space that separates the inputs based on the property
Findings  [^1] All results and experiments accessible on github:
  • We demonstrate the ability of diagnostic classifiers to effectively highlight known graph-theoretic and domain-specific properties in GNN learned latent representations (fig. 5).
  • We explore how different regularization techniques (none, $L_{2}$ weight decay, dropout) affect the representation of graph properties within the same GNN architecture (fig. 10).
  • We compare how various GNN architectures (GCN, R-GCN, GIN, GAT) differ in their ability to represent graph properties, analyzing whether these differences align with their mathematical frameworks (table 6).
  • We apply this pipeline to a toxicity dataset showing that probed graph properties align with chemical knowledge (table 8) before exploring the pipeline’s inferential power on fMRI datasets, uncovering structural properties that might not yet have been extensively studied (table 23).
@@ 2.1 Graph Neural Networks
Nowadays we have some theoretical understanding of the representational restrictions and capabilities of Graph Neural Networks (GNNs) with regard to the Weisfeiler-Lehman test (Akhondzadeh et al., 2023). We know that this cannot capture certain graph properties, such as connectivity or triangle-freeness (Franks et al., 2024; Kiefer, 2020; Kriege et al., 2018), due to its reliance on local structure. This constraint is also present in (message passing) GNNs.
Graph Convolutional Network (GCN) (Kipf & Welling, 2017) are GNNs where for a single layer, the node representation is computed as: $\bm{X}^{\prime}=\sigma\left(\tilde{\bm{D}}^{-1/2}\cdot\tilde{\bm{A}}\cdot\tilde{\bm{D}}^{-1/2}\cdot\bm{X}\cdot\bm{W}\right)$. We know that GNNs which rely solely on local information, like the GCN and its relational variant (R-GCN(Schlichtkrull et al., 2018), cannot compute important graph properties, such as girth and diameter or eigenvector centrality Garg et al. (2020). We are therefore also investigating more globally aware networks like GAT (Graph Attention Network) (Veličković et al., 2018) and GIN (Graph Isomorphism Network) (Xu et al., 2019).
GAT makes use of self-attention and is thereby more expressive than the GCN. However, its reliance on feature-dependent weights and structure-free normalisation limits its ability to capture specific structural properties that do not directly depend on edges. This is particularly true for tasks where node features alone are not enough, and global graph structures are crucial (e.g., tasks requiring knowledge of subgraphs or non-local patterns). GIN aggregates node features in a way that mimics the Weisfeiler-Lehman test for graph isomorphism, and with its strong inductive learning capabilities, it is likely to excel at encoding complex graph properties and solving classification tasks.
@@ 2.2 Graph properties
Graph theory is a branch of mathematics that studies the properties and relationships of graphs. Graphs can be undirected or directed and analysed through both local and global properties. Local properties like node degree which count the number of connections a node has, identifying highly connected hub nodes, or the clustering coefficient which measures how well a node’s neighbours are interconnected, capturing the local density of connections and giving the node a score, are based on a node with regard to its neighbour. In contrast, global properties such as diameter and characteristic path length assess the overall structure. They indicate how far nodes are from one another and how efficiently information can spread through the network. Global graph properties can be associated with higher level complex systems’ characteristics like the presence of some repeated motifs in the sub structures of the graphs or information-flow properties.
We can distinguish different global properties, basic ones like the number of nodes a graph has, clustering and centrality ones, graph motifs and substructures, spectral and small-world properties.
As an higher-order analysis, the recurrence of specific motifs within network substructures—such as triangles, cliques, or feed-forward loops can be seen as the fundamental building blocks that dictate the system’s functionality and resilience. Small-worldness  [^2] We are using the Small-World Index, $SWI=\left(\frac{L-L_{l}}{L_{r}-L_{l}}\right)\times\left(\frac{C-C_{r}}{C_{l}-C_{r}}\right)$ in our experiment because it provides a more balanced and robust measure of small-world properties. Unlike the Small-World Quotient: $Q=\frac{C/C_{r}}{L/L_{r}}$, which can be sensitive to network size and degree, $SWI$ normalises both the clustering coefficient and average path length with respect to both random and lattice reference graphs. This dual normalisation approach ensures that $SWI$ is less prone to false positives or negatives, making it a more reliable metric for our analysis Neal (2017). , as characterised by Barabási Albert & Barabási (2002), reveal how networks can maintain short path lengths despite their expansive size and sparse connectivity. This kind of higher order properties are very interesting in order to understand how the macroscopic behaviour of complex systems emerges from the intricate interplay of their microscopic components Barabási et al. (2002). For example how diseases spread in social networks, how neurons interact in the brain, or how information propagates through the Internet. GNNs synthesise local topological features into global structures, abstract these representations into higher-order graph attributes. Probing their learnt representations should act as a scalable proxy to investigate how global arrangement and connectivity patterns influence a system’s function. In other terms, by dissecting these learned embeddings, we can possibly delve into the intricate relationships between a network’s macroscopic arrangement and its emergent behaviours.
A Graph $G=(V,E)$, $V$ the set of vertices, $E$ the set of edges, can be analysed through both local and global properties. Local properties (like node degree or clustering coefficient) are based on the neighbors of a node.
@@
In contrast, global properties (such as diameter and characteristic path length) assess the overall graph structure. Global graph properties can be associated with higher level complex systems’ characteristics like the presence of repeated motifs in the graphs or information-flow properties. See the appendix B for a list of local and global properties used in our experiments.
GNNs synthesise local topological features into global structures and then abstract these representations into higher-order graph attributes. Probing their learnt representations should act as a scalable proxy to investigate how global arrangement and connectivity patterns influence a system’s function. Based on the message passing paradigm in GNNs, as layers progress, one would expect an increased abstraction in the selection of graph properties. Initially, local features like node degree dominate, but deeper layers progressively capture more global properties, such as connectivity patterns and centrality.
Through hierarchical pooling or readout mechanisms, GNNs can aggregate node embeddings into a single, global graph-level embedding. Graphs that share structural similarities or patterns of interaction among nodes are organised closely in the embedding space, allowing the model to differentiate between classes of graphs, such as those with and without long paths.
@@ 2.3 Probing classifiers
In prior work (Hupkes et al., 2018) probing classifiers have been used for linguistic properties. Here, we adapt them for graph features. Unlike unsupervised techniques such as Principal Component Analysis (PCA) or T-SNE, which are useful to visualise input data with regard to the embedding latent space, we adopt a supervised framework to quantitatively assess how specific properties are encoded within the embedding space of DNNs. Let $g:f_{l}(x)\mapsto\hat{z}$ represent a probing classifier, used to map the learned intermediate representations from the original model $f$ to a specific property $\hat{z}$. The choice of a linear classifier for $g$ is motivated primarily by its simplicity. If a linear probe performs well, it suggests the existence of a hyperplane in the representation space that separates the inputs based on their properties, indicating linear separability.
@@
Another advantage of a simple linear probe is avoiding the risk that a more complex classifier might infer features that are not actually used by the network itself Hupkes et al. (2018). While other non-linear probes have been explored in the literature Belinkov (2021), even studies showing improved performance with complex probes maintain the same logic: $\operatorname{Perf}(g,f_{1},\mathcal{D}_{O},\mathcal{D}_{P})>\operatorname{Perf}(g,f_{2},\mathcal{D}_{O},\mathcal{D}_{P})$ holds across representations $f_{1}(x)$ and $f_{2}(x)$ when evaluated by a consistent probe $g$. This consistency ensures valid comparison, underscoring that if a property can be predicted well by a simple probe, it is likely relevant to the primary classification task.
From an information-theoretic perspective, training the probing classifier $g$ can be viewed as estimating the mutual information between the learned representations $f_{l}(x)$ and the property $z$. This mutual information is denoted as $\mathrm{I}(\mathbf{z};\mathbf{h})$, where $\mathbf{z}$ refers to the property and $\mathbf{h}$ represents the intermediate representations Belinkov (2021).
@@
This supervised approach allows us to define hyperplanes or higher-dimensional decision boundaries that partition the embedding space according to the chosen graph property. The $R^{2}$ score serves as this information-theoretic measure indicating how well the hyperplane divides the inputs in the embedding space. A $R^{2}$ near 1 indicates that the embeddings are highly informative about $\hat{z}$, suggesting that the neural model has internalized this property in a linearly accessible manner.
@@
By defining specific properties that could divide the embedding space and assessing how well the corresponding hyperplanes make the embedding space linearly separable, we gain quantitative insights into the abstract features aggregated within the embeddings. This method moves beyond mere hypothesis generation based on clustering patterns observed through techniques like PCA, providing a rigorous framework for understanding how well the embedding space represents complex graph properties. It can also be thought as complementary from the T-SNE and PCA visualisation techniques, as it provides a quantitative measure of the separability of the embeddings based on hypothesised properties of interest.
The best illustration of this comes with fig. 3. We illustrate the evolution of the separability of graphs in the embeddings also in fig. 5 using a T-SNE visualisation and the corresponding separability with the properties thanks to the probing. This highlight the most interesting results of the paper, showing that the separability of house-only and grid-only graphs in the negative class (purple) match with the presence of the property number of triangles in the 5th layer of the GIN architecture.
3 Related Work
Existing post-hoc GNN explanations methods can be classified into two main categories: instance-level and model-level methods Barredo Arrieta et al. (2020). See Agarwal et al. (2023); Dai et al. (2022) for nice reviews on the subject. In the realm of instance based methods, gradient-based methods use the gradients of the output with respect to the input or intermediate features to measure the importance of each component of the graph. Decomposition-based methods try to decompose the input graph into smaller subgraphs or paths that can account for the output. Surrogate-based methods use a simpler, more interpretable model to approximate the behaviour of the original GNN and provide explanations based on the surrogate model. And finally Perturbation-based methods which perturb the input graph by removing or adding nodes, edges, or features, and observe the changes in the output to identify the influential components.
The most mainstream technique, GNNExplainer Ying et al. (2019) achieves explanation by removing redundant edges from an input graph instance, maximising the mutual information between the distribution of subgraphs and the GNN’s prediction. It is able to provide an explanation both in terms of a subgraph of the input instance to explain, and a feature mask indicating the subset of input node features which is most responsible for the GNN’s prediction.
For model-based techniques, few methods come to mind Saha et al. (2022); Azzolin et al. (2023); Vu & Thai (2020); Wang et al. (2023); Xuanyuan et al. (2023); Yuan et al. (2020); Zhang et al. (2021). The most mainstream method seems to be XGNN Yuan et al. (2020). The authors of XGNN investigate the possible input characteristics used by a GNN for graph classification. But they formulate the problem as a reinforcement learning problem and generate graph patterns iteratively. Such an iterative approach is often intractable for large graphs. Moreover, it does not allow for both node classification and graph classification explanations, nor does it allow for an investigation of the learning process through the different layers of the GNN.
@@ 4 Datasets
inspired by (Agarwal et al., 2023) is designed to evaluate the compositionality of Graph Neural Network (GNN). It features two concepts: a 3x3 grid and a house-shaped graph made of five nodes. The dataset consists of Barabási-Albert (BA) graphs (Barabási, 2009) with a normal distribution of the number of nodes. The negative class includes a BA graph connected to either a grid or a house, while the positive class contains a BA graph connected to both a grid and a house (see fig. 2). In order to ensure that the average number of nodes is the same between classes, the number of nodes is a uniformly distributed between 6 and 21 for the grid graphs, between 7 and 22 for the house graphs, and between 1 and 16 when both are present. During generation, we ensure no test set leakage by removing isomorphisms. On 2,000 graphs, we perform an 80/20 train/test split.
@@
Figure 2: Examples of the grid-house dataset. There are graphs with only a house and only a 3x3 grid, these are in one class. Graphs with both a house and a grid are in the other class.
@@
In our paradigm, we hope probing functional connectivity matrices (FC) matrices (Farahani et al., 2019) of neurological disorders (ND) could help explore the link between stuctural properties of the brain’s functional connectivity and neurological disorders such as Autism Spectrum Disorders (ASD) and Major Depressive Disorders (MDD). Similarly that probed graph properties in toxic molecules align with chemical knowledge.

### v2

@@ Abstract
Deep neural networks (DNNs) achieve state-of-the-art performance on many tasks, but this often requires increasingly larger model sizes, which in turn leads to more complex internal representations. Explainability techniques (XAI) have made remarkable progress in the interpretability of ML models. However, the non-relational nature of Graph neural networks (GNNs) make it difficult to reuse already existing XAI methods. While other works have focused on instance-based explanation methods for GNNs, very few have investigated model-based methods and, to our knowledge, none have tried to probe the embedding of the GNNs for well-known structural graph properties. In this paper we present a model agnostic explainability pipeline for GNNs employing diagnostic classifiers. This pipeline aims to probe and interpret the learned representations in GNNs across various architectures and datasets, refining our understanding and trust in these models.
@@ 1 Introduction
Graph Neural Networks (GNNs) are pivotal in harnessing non-Euclidean graph-structured data (Kipf & Welling, 2017) for tasks ranging from social network analysis to bioinformatics. Despite their success, the black-box nature of GNNs poses significant challenges as classical XAI methods can’t be directly applied on GNNs due to the lack of a regular structure (e.g. vertices can have different degrees). In this case, explaining a prediction means identifying important parts of the relational structure, or input features of nodes. An issue is that finding the explanation is itself a combinatorial problem, making XAI methods for GNN intractable (Longa et al., 2023a; Ying et al., 2019).
@@
Previous surveys (Agarwal et al., 2023; Dai et al., 2022) highlighted the lack of comprehensive, robust and model-agnostic explainability methods. We also identified (sec C) few model-level explainability methods. They focus on explaining the decision-making process at a high level, often by generating graph patterns or motifs that influence the model’s predictions but none allow for the interpretability of intermediate layers and none highlight the role of graph properties. One paper identified the room for probing classifier (Akhondzadeh et al., 2023), as developed for Natural Language Processing (Giulianelli et al., 2018), (Belinkov, 2021). Their research question was ”will the hidden representation contain information about the number of hydrogen atoms or the presence of aromatic rings?”. We aim to address a more fundamental and general question : Does the hidden representation encode information about the graph-theoretic properties ? We propose a simple, model-agnostic, probing classifier pipeline to interpret GNN embeddings by probing for encoded graph properties (see fig. 1) across various architectures and datasets. We investigate both local properties like betweennes centrality, as well as global properties like average path length. To our knowledge, this is the first work to explore this direction.
We demonstrate the ability of diagnostic classifiers to effectively highlight graph-theoretic properties in GNN learned latent representations (fig. 7). We explore how different regularization techniques (none, $L_{2}$ weight decay, dropout) affect the representation of graph properties (fig. 14). We compare how various GNN architectures (GCN, R-GCN, GIN, GAT) differ in their ability to represent graph properties, analyzing whether these differences align with their mathematical frameworks (table 6). The approach is validated through applications to toxicity and fMRI datasets, confirming the alignment of probed properties with domain knowledge (table 8), before exploring the pipeline’s inferential power, uncovering structural properties that might not yet have been extensively studied (table 23).
@@
Figure 1: Illustration of the probing pipeline. First, a GNN is trained on a specific task, such as classifying between smaller graphs (label 0) and larger graphs (label 1). Similar to how a CNN might organize images of similar shapes or textures into distinct regions in its feature space, the GNN embeddings might arrange graphs based on structural properties like their diameter. Next, we extract embeddings from the internal layers of the GNN. These embeddings are used to train the probing model—in this case, a binary classifier tasked with determining whether the embeddings encode predictive information about the graph’s diameter. If a linear probe has good performance ($R^{2}$ score) then there exists a hyperplane in the representation space that separates the inputs based on the property
@@ 2.1 Graph Neural Networks
Nowadays we have some theoretical understanding of the representational restrictions and capabilities of Graph Neural Networks (GNNs) with regard to the Weisfeiler-Lehman test (Akhondzadeh et al., 2023). We know that these cannot capture certain graph properties, such as connectivity or triangle-freeness (Franks et al., 2024; Kiefer, 2020; Kriege et al., 2018), due to their reliance on local structure. This constraint is also present in (message passing) GNNs.
Graph Convolutional Network (GCN) (Kipf & Welling, 2017) are GNNs where for a single layer, the node representation is computed as: $\bm{X}^{\prime}=\sigma\left(\tilde{\bm{D}}^{-1/2}\cdot\tilde{\bm{A}}\cdot\tilde{\bm{D}}^{-1/2}\cdot\bm{X}\cdot\bm{W}\right)$. We know that GNNs which rely solely on local information, like the GCN and its relational variant (R-GCN(Schlichtkrull et al., 2018), cannot compute important graph properties, such as girth and diameter or eigenvector centrality (Garg et al., 2020). We are therefore also investigating more globally aware networks like GAT (Graph Attention Network) (Veličković et al., 2018) and GIN (Graph Isomorphism Network) (Xu et al., 2019).
GAT makes use of self-attention and is thereby more expressive than the GCN. However, its reliance on feature-dependent weights and structure-free normalisation limits its ability to capture specific structural properties that do not directly depend on edges. This is particularly true for tasks where node features alone are not enough, and global graph structures are crucial (e.g., tasks requiring knowledge of subgraphs or non-local patterns). GIN aggregates node features in a way that mimics the Weisfeiler-Lehman test. By using the MLP equivalent to an Injective Update Function, GIN avoids oversimplifying the aggregation step, making it as expressive as the WL test. Thus, it is likely to excel at encoding complex graph properties and solving classification tasks.
@@ 2.2 Graph properties
Graph theory is a branch of mathematics that studies the properties and relationships of graphs. Graphs can be undirected or directed and analysed through both local and global properties. Local properties (like node degree or clustering coefficient) are based on the neighbors of a node. In contrast, global properties (such as diameter and characteristic path length) assess the overall graph structure. Global graph properties can be associated with higher level complex systems’ characteristics like the presence of repeated motifs in the graphs or information-flow properties. See the appendix B for a list of local and global properties used in our experiments.
We can distinguish different global properties, basic ones like the number of nodes a graph has, clustering and centrality ones, graph motifs and substructures, spectral and small-world properties. As an higher-order analysis, the recurrence of specific motifs within network substructures—such as triangles, cliques, or feed-forward loops can be seen as the fundamental building blocks that dictate the system’s functionality and resilience. Small-worldness, as characterised by Barabási (Albert & Barabási, 2002), reveal how networks can maintain short path lengths despite their expansive size and sparse connectivity. This kind of higher order properties are very interesting in order to understand how the macroscopic behaviour of complex systems emerges from the intricate interplay of their microscopic components (Barabási et al., 2002). For example how diseases spread in social networks, how neurons interact in the brain, or how information propagates through the Internet.
GNNs synthesise local topological features into global structures, abstract these representations into higher-order graph attributes. Each layer progressively expands the receptive field, mirroring how hierarchical feature learning works in convolutional neural networks (CNNs) for images. Probing their learnt representations should act as a scalable proxy to investigate how global arrangement and connectivity patterns influence a system’s function. In other terms, by dissecting these learned embeddings, we can possibly delve into the intricate relationships between a network’s macroscopic arrangement and its emergent behaviours. Based on the message passing paradigm in GNNs, as layers progress, one would expect an increased abstraction in the selection of graph properties. Initially, local features like node degree dominate, but deeper layers progressively capture more global properties, such as connectivity patterns and centrality. Through hierarchical pooling or readout mechanisms, GNNs can aggregate node embeddings into a single, global graph-level embedding. Graphs that share structural similarities or patterns of interaction among nodes are organised closely in the embedding space, allowing the model to differentiate between classes of graphs, such as those with and without long paths (fig. 1). This is why, AI engineers are likely to focus on the validation framework, emphasizing the alignment between probing results and the predictions derived from the message-passing paradigm—particularly insights revealed in the initial layers of the model. In contrast, from the standpoint of domain researchers—such as chemists or neuroscientists— the most compelling aspect is found in the later layers of the model, where the abstract representations become increasingly capable of rendering the problem linearly separable, thus facilitating clear interpretability of classification decisions and offering domain-specific interpretations.
@@ 2.3 Probing classifiers
In prior work (Hupkes et al., 2018) probing classifiers have been used for linguistic properties. Here, we adapt them for graph features. Unlike unsupervised techniques such as PCA or T-SNE, which are useful to visualize input data with regard to the embedding latent space, we adopt a supervised framework to quantitatively assess how specific properties are encoded within the embedding space of GNNs. Let $g:f_{l}(x)\mapsto\hat{z}$ represent a probing classifier, used to map the learned intermediate representations from the original model $f$ to a specific property $\hat{z}$. The choice of a linear classifier for $g$ is motivated primarily by its simplicity. If a linear probe performs well, it suggests the existence of a hyperplane in the representation space that separates the inputs based on their properties, indicating linear separability.
@@
Another advantage of a simple linear probe is avoiding the risk that a more complex classifier might infer features that are not actually used by the network itself (Hupkes et al., 2018). While other non-linear probes have been explored in the literature (Belinkov, 2021), even studies showing improved performance with complex probes maintain the same logic: $\operatorname{Perf}(g,f_{1},\mathcal{D}_{O},\mathcal{D}_{P})>\operatorname{Perf}(g,f_{2},\mathcal{D}_{O},\mathcal{D}_{P})$ holds across representations $f_{1}(x)$ and $f_{2}(x)$ when evaluated by a consistent probe $g$. This consistency ensures valid comparison, underscoring that if a property can be predicted well by a simple probe, it is likely relevant to the primary classification task. From an information-theoretic perspective, training the probing classifier $g$ can be viewed as estimating the mutual information between the learned representations $f_{l}(x)$ and the property $z$. This mutual information is denoted as $\mathrm{I}(\mathbf{z};\mathbf{h})$, where $\mathbf{z}$ refers to the property and $\mathbf{h}$ represents the intermediate representations (Belinkov, 2021).
@@
This supervised approach allows us to define hyperplanes or higher-dimensional decision boundaries that partition the embedding space according to the chosen graph property. The $R^{2}$ score serves as this information-theoretic measure indicating how well the hyperplane divides the inputs in the embedding space. The R2 score (, see appendix A for a formal definition) measures the proportion of variance in the dependent variable that is predictable from the independent variable(s). A $R^{2}$ near 1 indicates that the embeddings are highly informative about $\hat{z}$, suggesting that the neural model has internalized this property in a linearly accessible manner.
@@
By defining specific properties that could divide the embedding space and assessing how well the corresponding hyperplanes make the embedding space linearly separable, we gain quantitative insights into the abstract features aggregated within the embeddings. This method moves beyond mere hypothesis generation based on clustering patterns observed through techniques like PCA, providing a rigorous framework for understanding how well the embedding space represents complex graph properties. It can also be thought as complementary from the T-SNE and PCA visualisation techniques, as it provides a quantitative measure of the separability of the embeddings based on hypothesised properties of interest. We illustrate the evolution of the separability of graphs in the embeddings in figs. 4, 7 and 16 using a T-SNE visualisation and the corresponding separability with the properties thanks to the probing.
@@ 3 Datasets
inspired by (Agarwal et al., 2023) is designed to evaluate the compositionality of GNNs. It features two concepts: a 3x3 grid and a house-shaped graph made of five nodes. The dataset consists of Barabási-Albert (BA) graphs (Barabási, 2009) with a normal distribution of the number of nodes. The negative class includes a BA graph connected to either a grid or a house, while the positive class contains a BA graph connected to both a grid and a house (see fig. 2). In order to ensure that the average number of nodes is the same between classes, the number of nodes is a uniformly distributed between 6 and 21 for the grid graphs, between 7 and 22 for the house graphs, and between 1 and 16 when both are present. During generation, we ensure no test set leakage by removing isomorphisms. On 2,000 graphs, we perform an 80/20 train/test split.
@@
The Grid-House dataset serves a critical purpose in our study: it offers a controlled and well-defined environment for rigorously testing our hypotheses. The simplicity of these controlled constraints allow us to verify whether the GNN operates as intended in a setting where extraneous factors are minimized. If a model struggles with this dataset, it would likely underperform on real-world, more complex graphs, underscoring its diagnostic value.
Figure 2: Illustration of the grid-house dataset. The first class (0) include graphs with either a house (square+triangle) either a 3x3 grid (4 squares). The second class include both a house and a grid.